OpenStudy (anonymous):

What is the 9th term of this arithmetic sequence? 2a - b, a, b, .....

OpenStudy (turingtest):

let's find the common difference between each term I guess...

OpenStudy (turingtest):

the common difference is d=a-(2a-b)=b-a so we have up to the third term b 4th term b+(b-a)=2b-a 5th term 2b-a+(b-a)=3b-2a with the common difference you should be able to figure out what the nth term is. if not, just do it manually ;)

OpenStudy (anonymous):

Thanks!!

OpenStudy (turingtest):

welcome :)